# Multistep Consistency Models

### v2

@@ Abstract
Multistep Consistency Models work really well in practice. By increasing the sample budget from a single step to 2-8 steps, we can train models more easily that generate higher quality samples, while retaining much of the sampling speed benefits. Notable results are 1.4 FID on Imagenet 64 in 8 step and 2.1 FID on Imagenet128 in 8 steps with consistency distillation, using simple losses without adversarial training. We also show that our method scales to a text-to-image diffusion model, generating samples that are close to the quality of the original model.
@@ 2 Background: Diffusion Models
In contrast, consistency models (Song et al., 2023; Song & Dhariwal, 2023) aim to learn a direct mapping from noise to data. Consistency models are constrained to predict ${\bm{x}}=f({\bm{z}}_{0},0)$, and are further trained by learning to be consistent, minimizing:
@@
An important hyperparameter in consistency models is the gap between the model evaluations at $t$ and $s$. For CT large gaps incurs a bias, but the solutions are propagated through diffusion time more quickly. On the other hand, when $s\to t$ the bias tends to zero but it takes much longer to propagate information through diffusion time. In practice a step schedule $N(\cdot)$ is used to anneal the step size $t-s=1/N(\cdot)$ over the course of training.
@@
Another helpful equations is the inverse of DDIM (Salimans & Ho, 2022), originally proposed to find a natural way parameterize a student diffusion model when a teacher defines the sampling procedure in terms of ${\bm{z}}_{t}$ to ${\bm{z}}_{s}$. The equation takes in ${\bm{z}}_{t}$ and ${\bm{z}}_{s}$, and produces ${\bm{x}}$ for which $\operatorname{DDIM}_{t\to s}({\bm{x}},{\bm{z}}_{t})={\bm{z}}_{s}$. It can be derived by rearranging terms from the $\operatorname{DDIM}$ equation:
@@ 3 Multistep Consistency Models
  $T_{\mathrm{step}}=\operatorname{round}(N(i)/\mathrm{steps})$
  $\mathrm{step}\sim\mathcal{U}(0,\mathrm{steps}-1)$, $n_{rel}\sim\mathcal{U}(1,T_{\mathrm{step}})$
  $t_{\mathrm{step}}=\mathrm{step}/\mathrm{steps}$
@@
  $\hat{{\bm{z}}}_{t_{\mathrm{step}}}=\operatorname{DDIM}_{t\to t_{\mathrm{step}}}(\hat{{\bm{x}}},{\bm{z}}_{t})$
@@
3.1 General description
Multistep consistency split up diffusion time into equal segments to simplify the modelling task. Recall that a consistency model must learn to integrate the full ODE integral. This mapping can become very sharp and difficult to learn when it jumps between modes of the target distribution as can be seen in Figure 1. A consistency loss can be seen as an objective that aims to approximate a path integral by minimizing pairwise discrepancies. Multistep consistency generalizes this approach by breaking up the integral into multiple segments. Originally, consistency runs until time-step 00, evaluated at some time $t>0$. A consistency model should now learn to integrate the DDIM path until 00 and predict the corresponding ${\bm{x}}$. Instead, we can generalize the consistency loss to targets $z_{t_{\mathrm{step}}}$ instead of ${\bm{x}}$ ($\approx{\bm{z}}_{0}$). It turns out that the DDIM equation can be used to operate on ${\bm{z}}_{t_{\mathrm{step}}}$ for different times $t_{\mathrm{step}}$, which allows us to express the multi-step consistency loss as:
@@ 3.1 General description
where $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}=\operatorname{DDIM}_{s\to t_{\mathrm{step}}}(\operatorname{nograd}f({\bm{z}}_{s},s))$ and where the teaching step ${\bm{z}}_{s}=\operatorname{aDDIM}_{t\to s}(x,{\bm{z}}_{t})$. For now it suffices to think $\operatorname{aDDIM}$ as $\operatorname{DDIM}$. It will be described in detail in section 3.2. In fact, one can drop-in any deterministic sampler (or integrator) in place of $\operatorname{aDDIM}$ in the case of distillation.
A model can be trained on directly on this loss in $z$ space, however make the loss more interpretable and relate it more closely to standard diffusion, we re-parametrize the loss to $x$-space using:
@@
This allows the usage of existing losses from diffusion literature, where we have opted for $v$-loss (equivalent to $\mathrm{SNR}+1$ weighting) because of its prior success in distillation (Salimans & Ho, 2022).
As noted in (Song et al., 2023), consistency in itself is not sufficient to distill a path (always predicting 00 is consistent) and one needs to ensure that the model cannot collapse to these degenerate solutions. Indeed, in our specification observe that that $\operatorname{DDIM}_{t_{\mathrm{step}}\to t_{\mathrm{step}}}({\bm{z}}_{t_{\mathrm{step}}},\hat{\bm{x}})={\bm{z}}_{t_{{\mathrm{step}}}}$ and that for the final step in each segment $s=t_{\mathrm{step}}$. As such, the loss of the final step cannot be degenerate and is equal to:
$\small||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t_{\mathrm{step}}\to t}({\bm{z}}_{t_{\mathrm{step}}})||.$ (8)
@@
Consistency training learns to integrate the probability flow through time, whereas standard diffusion models learn a path guided by an expectation $\hat{{\bm{x}}}=\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$ that necessarily has to change over time for non-trivial distributions. There are two simple reasons that for many steps, Multistep CT converges to a diffusion model. 1) At the beginning of a step (specifically $t={t_{\mathrm{step}}}+\frac{1}{T}$) the objectives are identical. Secondly, 2) when the number of steps equals the total number of steps $T$, then every step is equal to the diffusion objective. This can be observed by studying Algorithm 1: let $t=t_{\mathrm{step}}+1/T$. For consistency training, aDDIM reduces to DDIM and observe that in this case $s=t_{\mathrm{step}}$. Hence, under a well-defined model $f$ (such as a $v$-prediction one) $\operatorname{DDIM}_{s\to t_{\mathrm{step}}}$ does nothing and simply produces $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}={\bm{z}}_{s}$. Also observe that $\hat{{\bm{z}}}_{t_{\mathrm{step}}}=\hat{{\bm{z}}}_{s}$. Further simplification yields:
$w(t)||{\bm{x}}_{\mathrm{diff}}||=\operatorname{invDDIM}_{t\to s}(\hat{{\bm{z}}}_{s},{\bm{z}}_{t})-\hat{{\bm{x}}}=w(t)||{\bm{x}}-\hat{{\bm{x}}}||$ (9)
Where $||{\bm{x}}-\hat{{\bm{x}}}||$ is the distance between the true datapoint and the model prediction weighted by $w(t)$, which is typical for standard diffusion. Interestingly, in (Song & Dhariwal, 2023) it was found that Euclidean ($\ell_{2}$) distances typically work better than for consistency models than the more usual squared Euclidean distances ($\ell_{2}$ squared). We follow their approach because it tended to work better especially for smaller steps, which is a deviation from standard diffusion. Because multistep consistency models tend towards diffusion models, we can state two important hypotheses:
@@
    As the number of steps increases, Multistep CMs will rival diffusion model performance, giving a direct trade-off between sample quality and duration.
@@
  Sample ${\bm{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
  for $t$ in $(\frac{\mathrm{steps}}{\mathrm{steps}},\ldots,\frac{1}{\mathrm{steps}})$ where $s=t-\frac{1}{\mathrm{steps}}$ do
@@
3.2 The Adjusted DDIM (aDDIM) sampler.
Popular methods for distilling diffusion models, including the method we propose here, rely on deterministic sampling through numerical integration of the probability flow ODE. In practice, numerical integration of this ODE in a finite number of steps incurs error. For the DDIM integrator (Song et al., 2021a) used for distilling diffusion models in both consistency distillation (Song et al., 2023) and progressive distillation (Salimans & Ho, 2022; Meng et al., 2022) this integration error causes samples to become blurry. To see this quantitatively, consider a hypothetical perfect sampler that first samples ${\bm{x}}^{*}\sim p({\bm{x}}|{\bm{z}}_{t})$, and then samples ${\bm{z}}_{s}$ using
@@
$\small||{\bm{z}}^{*}_{s}||^{2}-||{\bm{z}}^{\text{DDIM}}_{s}||^{2}=\text{trace}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]),$ (12)
@@
The norm of the DDIM iterates is thus too small, reflecting the lack of noise addition in the sampling algorithm. Alternatively, we could say that the model prediction $\hat{{\bm{x}}}\approx\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$ is too smooth. Second-order integrators like the Heun sampler proposed by Karras et al. (2022) aim to remove this bias by using the prediction at $s$ in addition to that at $t$, but in practice these methods did not significantly improve sample quality in our experiments.
Figure 3: Comparison of the vanilla and adjusted DDIM sampler for the small ImageNet 128 diffusion model without distillation.
Currently, the best sample quality is achieved with stochastic samplers, which can be tuned to add exactly enough noise to undo the oversmoothing caused by numerical integration. However, current distillation methods are not well suited to distilling these stochastic samplers directly. Here we therefore propose a new deterministic sampler that aims to achieve the norm increasing effect of noise addition in a deterministic way. It turns out we can do this by making a simple adjustment to the DDIM sampler, and we therefore call our new method Adjusted DDIM (aDDIM). Our modification is heuristic and is not more theoretically justified than the original DDIM sampler. However, empirically we find aDDIM to work very well leading to improved FID scores (Fig. 3) and thus a stronger deterministic teacher.
@@ 4 Related Work
Multistep Consistency Models are a direct combination of (Song et al., 2023; Song & Dhariwal, 2023) and TRACT (Berthelot et al., 2023). Compared to consistency models, we propose to operate on multiple stages, which simplifies the modelling task and improves performance significantly. On the other hand, TRACT limits itself to distillation and uses the self-evaluation from consistency models to distill models over multiple stages. The stages are progressively reduced to either one or two stages and thus steps. The end-goal of TRACT is again to sample in either one or two steps, whereas we believe better results can be obtained by optimizing for a slightly larger number of steps. We show that this more conservative target, in combination with our improved sampler and annealed schedule, leads to significant improvements in terms of image quality that closes the gap between sample quality of standard diffusion and low-step diffusion-inspired approaches.
Earlier, DDIM (Song et al., 2021a) showed that deterministic samplers degrade more gracefully than the stochastic sampler used by Ho et al. (2020) when limiting the number of sampling steps. Karras et al. (2022) proposed a second order Heun sampler to reduce the number of steps (and function evaluations), while Jolicoeur-Martineau et al. (2021) studied different SDE integrators to reduce function evaluations. Progressive Distillation (Salimans & Ho, 2022; Meng et al., 2022) distills diffusion models in stages, which limits the number of model evaluations during training while exponentially reducing the required number of sampling steps with the number stages.
@@
Distillation to a few steps was very difficult to do using only simple distance metrics. Therefore, many works resort to a form of adversarial training. For example Luo et al. (2023) distill the knowledge from the diffusion model into a single-step model and Zheng et al. (2023) use specialized architectures to distill the ODE trajectory from a pre-created noise-sample pair dataset. A very similar approach to ours is Consistency Trajectory Models (CTMs) (Kim et al., 2023), which are trained to arbitrarily integrate to a given timestep. This is implemented by modifying the inputs of the denoising network to include an endpoint of the integration. Although CTMs produce very high quality image samples in few steps, their performance relies on adversarial training: Without it, CTMs cannot produce great samples and have a considerable gap in FID score. In contrast, our Multistep CMs can be trained with simple distance metrics and still achieve very good FID scores under few sampling steps. A possible explanation is that it is much easier to learn a handful of fixed integration trajectories (Multistep CMs) instead of every possible integration with arbitrary endpoints (CTMs). Another advantage of Multistep CMs is that the inputs to the denoising network are not changed, making fine-tuning of existing diffusion models to Multistep CMs very straightforward.
@@
Table 3: Comparison between PD (Salimans & Ho, 2022) and CT/CD on ImageNet64 on the small model.
@@ 5 Experiments
5.1 Quantitative Evaluation on ImageNet

### v3

@@ Abstract
Multistep Consistency Models work really well in practice. By increasing the sample budget from a single step to 2-8 steps, we can train models more easily that generate higher quality samples, while retaining much of the sampling speed benefits. Notable results are 1.4 FID on Imagenet 64 in 8 sampling steps and 2.1 FID on Imagenet128 in 8 sampling steps with consistency distillation, using simple losses without adversarial training. We also show that our method scales to a text-to-image diffusion model, generating samples that are close to the quality of the original model.
@@ 2 Background: Diffusion Models
In contrast, consistency models (Song et al., 2023; Song and Dhariwal, 2023) aim to learn a direct mapping from noise to data. Consistency models are constrained to predict ${\bm{x}}=f({\bm{z}}_{0},0)$, and are further trained by learning to be consistent, minimizing:
@@
An important hyperparameter in consistency models is the gap between the model evaluations at $t$ and $s$. For CT large gaps result in a bias, but the solutions are propagated through diffusion time more quickly. On the other hand, when $s\to t$ the bias tends to zero but it takes much longer to propagate information through diffusion time. In practice a step schedule $N(\cdot)$ is used to anneal the step size $t-s=1/N(\cdot)$ over the course of training.
@@
Another helpful equations is the inverse of DDIM (Salimans and Ho, 2022), originally proposed to find a natural way parameterize a student diffusion model when a teacher defines the sampling procedure in terms of ${\bm{z}}_{t}$ to ${\bm{z}}_{s}$. The equation takes in ${\bm{z}}_{t}$ and ${\bm{z}}_{s}$, and produces ${\bm{x}}$ for which $\operatorname{DDIM}_{t\to s}({\bm{x}},{\bm{z}}_{t})={\bm{z}}_{s}$. It can be derived by rearranging terms from the $\operatorname{DDIM}$ equation:
@@ 3.1 General description
  $N_{\mathrm{per\,segment}}=\!\operatorname{round}(N_{\mathrm{teacher}}(i)/\mathrm{student\_steps})$
  $\mathrm{step}\sim\mathcal{U}(0,\mathrm{student\_steps}-1)$
  $n_{rel}\sim\mathcal{U}(1,N_{\mathrm{per\,segment}})$
  $t_{\mathrm{step}}=\mathrm{step}/\mathrm{student\_steps}$
@@
Multistep consistency splits up diffusion time into equal segments to simplify the modelling task. Recall that a consistency model must learn to integrate the full ODE integral. This mapping can become very sharp and difficult to learn when it jumps between modes of the target distribution as can be seen in Figure 1. A consistency loss can be seen as an objective that aims to approximate a path integral by minimizing pairwise discrepancies. Multistep consistency generalizes this approach by breaking up the integral into multiple segments. Originally, consistency runs until time-step 00, evaluated at some time $t>0$. A consistency model should now learn to integrate the DDIM path until 00 and predict the corresponding ${\bm{x}}$. Instead, we can generalize the consistency loss to targets $z_{t_{\mathrm{step}}}$ instead of ${\bm{x}}$ ($\approx{\bm{z}}_{0}$). It turns out that the DDIM equation can be used to operate on ${\bm{z}}_{t_{\mathrm{step}}}$ for different times $t_{\mathrm{step}}$, which allows us to express the multi-step consistency loss as:
@@
where $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}=\operatorname{DDIM}_{s\to t_{\mathrm{step}}}(\operatorname{nograd}f({\bm{z}}_{s},s),{\bm{z}}_{s})$ and where the teaching step ${\bm{z}}_{s}=\operatorname{aDDIM}_{t\to s}(x,{\bm{z}}_{t})$ is an approximation of the probability flow ODE. For now it suffices to think of $\operatorname{aDDIM}$ as $\operatorname{DDIM}$. It will be described in detail in section 3.2. In fact, one can drop-in any deterministic sampler (or integrator) in place of $\operatorname{aDDIM}$ in the case of distillation.
A model can be trained directly on this loss in $z$ space, however make the loss more interpretable and relate it more closely to standard diffusion, we re-parametrize the loss to $x$-space using:
@@
This allows the usage of existing losses from diffusion literature, where we have opted for $v$-loss (equivalent to $\mathrm{SNR}+1$ weighting) because of its prior success in distillation (Salimans and Ho, 2022).
As noted in (Song et al., 2023), consistency in itself is not sufficient to distill a path (always predicting 00 is consistent) and one needs to ensure that the model cannot collapse to these degenerate solutions. Indeed, in our specification observe that when $s=t_{\mathrm{step}}$ then $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}=\operatorname{DDIM}_{s\to t_{\mathrm{step}}}({\bm{z}}_{s},\hat{\bm{x}})={\bm{z}}_{s}$. As such, the loss of the final step cannot be degenerate because it is equal to:
$\small||f({\bm{z}}_{t},t)-\operatorname{invDDIM}_{t\to s}({\bm{z}}_{s},{\bm{z}}_{t})||.$ (8)
@@
Consistency training learns to integrate the probability flow through time, whereas standard diffusion models learn a path guided by an expectation $\hat{{\bm{x}}}=\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$ that necessarily has to change over time for non-trivial distributions. There are two simple reasons that for many student steps, Multistep CT converges to a diffusion model. 1) At the beginning of a step (specifically $t={t_{\mathrm{step}}}+\frac{1}{T}$) the objectives are identical. Secondly, 2) when the number of student steps equals the number of teacher steps $T$, then every step is equal to the diffusion objective. This can be observed by studying Algorithm 1: let $t=t_{\mathrm{step}}+1/T$. For consistency training, aDDIM reduces to DDIM and observe that in this case $s=t_{\mathrm{step}}$. Hence, under a well-defined model $f$ (such as a $v$-prediction one) $\operatorname{DDIM}_{s\to t_{\mathrm{step}}}$ does nothing and simply produces $\hat{{\bm{z}}}_{\mathrm{ref},t_{\mathrm{step}}}={\bm{z}}_{s}$. Also observe that $\hat{{\bm{z}}}_{t_{\mathrm{step}}}=\hat{{\bm{z}}}_{s}$. Further simplification yields:
$w(t)||{\bm{x}}_{\mathrm{diff}}||=w(t)||\operatorname{invDDIM}_{t\to s}({\bm{z}}_{s},{\bm{z}}_{t})-\hat{{\bm{x}}}||=w(t)||{\bm{x}}-\hat{{\bm{x}}}||$ (9)
Where $||{\bm{x}}-\hat{{\bm{x}}}||$ is the distance between the true datapoint and the model prediction weighted by $w(t)$, which is typical for standard diffusion. Interestingly, in (Song and Dhariwal, 2023) it was found that Euclidean ($\ell_{2}$) distances typically work better than for consistency models than the more usual squared Euclidean distances ($\ell_{2}$ squared). We followed their approach because it tended to work better especially for smaller number of student steps, which is a deviation from standard diffusion. Because multistep consistency models tend towards diffusion models, we can state two important hypotheses:
@@
    As the number of student steps increases, Multistep CMs will rival diffusion model performance, giving a direct trade-off between sample quality and duration.
Note that this trade-off requires training a new Multistep CM for each of the desired student steps, but given that one starts from a pretrained model, one expects that finetuning requires a fraction of the original training budget.
@@
  Sample ${\bm{z}}_{1}\sim\mathcal{N}(0,{\mathbf{I}})$, $T=\mathrm{student\_steps}$
  for $t$ in $(\frac{T}{T},\ldots,\frac{1}{T})$ where $s=t-\frac{1}{T}$ do
@@
Popular methods for distilling diffusion models, including the method we propose here, rely on deterministic sampling through numerical integration of the probability flow ODE. In practice, numerical integration of this ODE in a finite number of teacher steps incurs error. For the DDIM integrator (Song et al., 2021a) used for distilling diffusion models in both consistency distillation (Song et al., 2023) and progressive distillation (Salimans and Ho, 2022; Meng et al., 2022) this integration error causes samples to become blurry. To see this quantitatively, consider a hypothetical perfect sampler that first samples ${\bm{x}}^{*}\sim p({\bm{x}}|{\bm{z}}_{t})$, and then samples ${\bm{z}}_{s}$ using
@@
$\small\mathbb{E}\big{[}||{\bm{z}}^{*}_{s}||^{2}-||{\bm{z}}^{\text{DDIM}}_{s}||^{2}\big{|}{\bm{z}}_{t}\big{]}=\text{trace}(\mathrm{Var}[{\bm{z}}_{s}|{\bm{z}}_{t}]),$ (12)
@@
The norm of the DDIM iterates is thus too small, reflecting the lack of noise addition in the sampling algorithm. Alternatively, we could say that the model prediction $\hat{{\bm{x}}}\approx\mathbb{E}[{\bm{x}}|{\bm{z}}_{t}]$ is too smooth.
Figure 3: Comparison of sampling methods for the small ImageNet 64 (left) and ImageNet 128 (right) models without distillation. The Heun (with sampler adds a second order correction to DDIM.
[…]
Here we therefore propose a new deterministic sampler that aims to achieve the norm increasing effect of noise addition in a deterministic way, with a single evaluation. It turns out we can do this by making a simple adjustment to the DDIM sampler, and we therefore call our new method Adjusted DDIM (aDDIM). Our modification is heuristic and is not more theoretically justified than the original DDIM sampler. However, empirically we find aDDIM to work very well leading to improved FID scores (Fig. 3) and thus a stronger deterministic teacher.
[…]
@@
Note that aDDIM only replaced the teacher steps. The student model uses a vanilla DDIM step which learns to predict the trajectory of the teacher with aDDIM. The student DDIM step only serves as a convenient output parameterization, and the student could just as well predict $z_{s}$ directly.
@@ 4 Related Work
Multistep Consistency Models are a direct combination of (Song et al., 2023; Song and Dhariwal, 2023) and TRACT (Berthelot et al., 2023). Compared to consistency models, we propose to operate on multiple stages, which simplifies the modelling task and improves performance significantly. On the other hand, TRACT limits itself to distillation and uses the self-evaluation from consistency models to distill models over multiple stages. The stages are progressively reduced to either one or two stages and thus steps. The end-goal of TRACT is again to sample in either one or two steps, whereas we believe better results can be obtained by optimizing for a slightly larger number of steps. We show that this more conservative target, in combination with our improved sampler and annealed schedule, leads to significant improvements in terms of image quality that closes the gap between sample quality of standard diffusion and low-step diffusion-inspired approaches.
Earlier, DDIM (Song et al., 2021a) showed that deterministic samplers degrade more gracefully than the stochastic sampler used by Ho et al. (2020) when limiting the number of sampling steps. Karras et al. (2022) proposed a second order Heun sampler to reduce the number of steps (and function evaluations), while Jolicoeur-Martineau et al. (2021) studied different SDE integrators to reduce function evaluations. Progressive Distillation (Salimans and Ho, 2022; Meng et al., 2022) distills diffusion models in stages, which limits the number of model evaluations during training while exponentially reducing the required number of sampling steps with the number stages.
@@
Low-step distillation typically suffers from degraded sample quality. Therefore, many works resort to a form of adversarial training. For example Luo et al. (2023) distill the knowledge from the diffusion model into a single-step model and Zheng et al. (2023) use specialized architectures to distill the ODE trajectory from a pre-created noise-sample pair dataset. Concurrent to our work (although released a few months earlier), Consistency Trajectory Models (CTMs) (Kim et al., 2023) also generalize CMs go multiple steps, and are trained to arbitrarily integrate to a given timestep. This is implemented by modifying the inputs of the denoising network to include an endpoint of the integration. Although CTMs produce very high quality image samples in a few steps, their performance relies on adversarial training: Without it, CTMs cannot produce great samples and have a considerable gap in FID score. In contrast, our simpler Multistep CMs can be trained with distance metrics and still achieve very good FID scores under a few sampling steps. A possible explanation is that it is much easier to learn a handful of fixed integration trajectories (Multistep CMs) instead of every possible integration with arbitrary endpoints (CTMs). Another advantage of Multistep CMs is that the inputs to the denoising network are not changed, making fine-tuning of existing diffusion models to Multistep CMs very straightforward. Finally, adversarial training comes with its own problems such as mode collapse and training instabilities.
@@
Table 3: Comparison between PD (Salimans and Ho, 2022) and CT/CD on ImageNet64 on the small model.
@@ 5 Experiments
5.1 Evaluation on ImageNet
Table 4: Ablation of the aDDIM teacher on ImageNet64.
Student Steps DDIM aDDIM
1 3.91 4.35
2 1.99 2.02
4 1.77 1.68
8 1.70 1.58
16 1.72 1.54
For our ImageNet experiments we trained diffusion models on ImageNet64 and ImageNet128 in a base and large variant. We initialize the consistency models from the pre-trained diffusion model weights which we found to greatly increase robustness and convergence. Both consistency training and distillation are used. Classifier Free Guidance (Ho and Salimans, 2022) was used only on the base ImageNet128 experiments. For all other experiments we did not use guidance because it did not significantly improve the FID scores of the diffusion model. All consistency models are trained for $200,000$ steps with a batch size of $2048$ and a teacher step schedule that anneals from $64$ to $1280$ in $100.000$ train steps with an exponential schedule.
In Table 1 the performance improves when the student step count increases. There are generally two patterns we observe: As the student steps increase, performance improves. This validates our hypothesis that more student steps are a useful trade-off between sample quality and speed. Conveniently, this happens very early: even on a complicated dataset such as ImageNet128, our base model variant is able to achieve 2.1 FID with just $8$ student steps.